If 10 people can finish a job in 30 days, In how many the job can be completed by 30 persons?
Answer by checkley79(3341)

10/30=30/X CROSS MULTIPLY
30X=30*10
30X=300
X=300/30
X=10 DAYS TO COMPLETE THE JOB WITH 30 WORKERS.